### v1

# Generative Physical AI in Vision: A Survey

Abstract
Generative Artificial Intelligence (AI) has rapidly advanced the field of computer vision by enabling machines to create and interpret visual data with unprecedented sophistication. This transformation builds upon a foundation of generative models to produce realistic images, videos, and 3D/4D content. Traditionally, generative models primarily focus on visual fidelity while often neglecting the physical plausibility of generated content. This gap limits their effectiveness in applications requiring adherence to real-world physical laws, such as robotics, autonomous systems, and scientific simulations. As generative AI evolves to increasingly integrate physical realism and dynamic simulation, its potential to function as a “world simulator" expands—enabling the modeling of interactions governed by physics and bridging the divide between virtual and physical realities. This survey systematically reviews this emerging field of physics-aware generative AI in computer vision, categorizing methods based on how they incorporate physical knowledge—either through explicit simulation or implicit learning. We analyze key paradigms, discuss evaluation protocols, and identify future research directions. By offering a comprehensive overview, this survey aims to help future developments in physically grounded generation for vision. The reviewed papers are summarized at https://github.com/BestJunYu/Awesome-Physics-aware-Generation.
1 Introduction
Generative learning has long been a foundational pillar of modern computer vision, addressing key challenges in understanding, synthesizing, and manipulating visual data. Over the past decade, this field has witnessed rapid evolution of diverse generative models, including Variational Autoencoders (VAEs)[1, 2], Generative Adversarial Networks (GANs)[3], Diffusion Models (DMs)[4, 5, 6], Neural Radiance Fields (NeRFs)[7], Gaussian Splatting (GS)[8], and Visual Autoregressive Models (VARs)[9]. These models have continually pushed the boundaries of generative learning, leveraging increasingly powerful architectures to capture the underlying distributions of visual data. The goal is to enable machines to reason about the visual world in ways that mirror human creativity and understanding, by imagining novel instances of visual content under unseen scenarios. Among these advancements, diffusion models have emerged as particularly noteworthy for their ability to produce highly realistic outputs. By iteratively refining random noise through learned denoising processes, diffusion models demonstrate exceptional robustness and versatility, making them a cornerstone of recent generative methodologies. The applications of generative models span diverse modalities of visual contents, including image generation with semantics understanding, video generation with dynamic temporal understanding, 3D content generation with enhanced spatial understanding[10, 11, 12], and 4D contents with more complex and combined understanding[13, 14, 15, 16, 17, 18, 19]. These advancements underscore the growing potential of generative learning in the broad landscape of increasingly sophisticated visual tasks.
Figure 1: The picture of generative AI in vision and the role of physics-aware generation on the path towards building a generative world simulator.
Among these different visual modalities, video generation has recently garnered significant attention within the field of generative learning, which serves as a much more challenging testbed for scaling up large generative models to handle higher dimensional data. This complexity stems not only from the spatial intricacies of individual frames but also from the temporal coherence required across sequences. Numerous commercial video generation models have been developed and have captured widespread public interest, such as OpenAI’s Sora[20], Google’s Veo2[21], Tencent’s Hunyuan[22], and Kuaishou’s Kling[23]. Video generation has been intensively studied in a range of formulations and settings, spanning from the vanilla unconditional generation[24, 25], to image-to-video generation[26, 27, 28, 29, 30, 31, 32, 33], text-to-video generation[24, 26, 34, 35, 25, 29, 36, 30, 37, 30], video-to-video generation[38, 39], and video editing or customization[40, 41, 42, 43]. Each of these settings addresses unique challenges, from preserving temporal continuity to incorporating semantic guidance from textual or visual inputs.
PS PU G PUG PAG-E PAG-I
Inputs (Observations) $\circ$ $\circ$ $\circ$ $\circ$
Inputs (Physics) $\circ$ $\circ$ $\circ$ $\circ$
Outputs (Observations) $\circ$
Outputs (Physics) $\circ$ $\circ$ $\circ$ $\circ$
Explicit Physical Model $\circ$
Understanding Physical World $\circ$ $\circ$
TABLE I: Comparison of Concepts. ✓: Yes. ✗: No. $\circ$: Optional.
PS: Physical simulation, PU: Physical understanding. G: Generation. PUG: Physics-unaware generation. PAG-E: Physics-aware generation with explicit physical simulation. PAG-I: Physics-aware generation without explicit physical simulation.
More importantly, video occupies a pivotal role in the future of generative AI in vision. The vast amount of video data available online encapsulates a rich repository of information about the real world, making video a medium through which generative AI can learn to model complex real-world phenomena. In this context, video can be regarded as a “language" of real-world decision making, offering the potential to bridge the digital and physical domains[44]. Video generation is posed to provide a unified interface as a “world model”[45] processing physical knowledge, analogous to how textual large language models (LLMs) process abstract knowledge. Such a model could facilitate a large number of downstream tasks, including autonomous driving, scientific simulation, robotics[46, 47, 48, 49, 50], and other forms of embodied intelligence. To realize this potential, the generation process should be capable of interacting with external controls from humans or other systems. This interactivity facilitates dynamic decision-making and the optimization of outcomes based on interactions, giving rise to what can be described as generative interactive environments[51, 44, 52, 53]. Video generation has been incorporated with various interactive control signals, such as motion vectors or trajectories[54, 55, 56, 57, 58], hand masks[59], latent actions[53, 60], robot operations[47], camera motions[61], demonstrations[62], and natural language descriptions[63, 64, 65]. These interactive elements highlight the versatility and adaptability of generative video models, paving the way for their evolution into world models.
However, a critical gap remains in the transition from generation to robust world modeling: the ability to faithfully understand and replicate real-world physics[66] (Fig. 1). Current state-of-the-art models are primarily optimized for visual realism in the pixel space, rather than the physical plausibility in the entity or concept space. For generative models to serve as simulators of the physical world, they must incorporate a deep understanding of physical laws, such as dynamics, causality, and material properties. This physics awareness is essential for going beyond generating visually appealing outputs to ensure that the content also aligns with the constraints and behaviors of the physical world.
Therefore, we provide this survey as a timely and comprehensive review of existing literature that attempts to embed physics-awareness into generative models. By examining these efforts, we hope to highlight the progress made so far, provide a clear structure of paradigms, and identify potential future directions.
Survey Scope. The scope of this survey is about generative models in computer vision that enhance the physical awareness of the generation outputs. Therefore, we exclude from our scope the literature that incorporates physical principles as prior knowledge or inductive biases in model or neural architecture design, such as Physics-Informed Neural Networks (PINNs)[67, 68], even if the task is related to generative learning, e.g., [69, 70, 71]. We focus on generative tasks, and thus exclude image processing tasks such as de-blurring, de-hazing, and enhancement from our scope, though we notice that a large body of these works incorporate physics. To concentrate on computer vision, pure graphic and rendering research combined with physical simulation is also excluded from our scope.
Comparison to Other Surveys. As delineated in our scope, this survey is orthogonal to existing surveys on physics-informed machine learning[72], physics-informed computer vision[73], and physics-informed artificial intelligence[74], since they emphasize the model design aspect with physical prior knowledge. Our survey focuses on generation with physical awareness, thus being distinct from existing surveys on generative models[75], diffusion models[76, 77], video diffusion models[78], diffusion-based video editing[79] in a general sense. Our survey also has a different scope compared to those in specific domains, such as human video or motion generation[80, 81, 82].
2 Formulation
In this section, we first provide a definition of physics-aware generation and related concepts, such as physical simulation and physical understanding. Based on the definition, we further identify the common paradigms of incorporating physics into generative models in vision, to provide a structural perspective for later sections of this survey.
2.1 Definitions
Let $P_{\theta}$ denotes a physical simulation model with physical parameters $\theta$, and $G$ represents a generative model, we can then provide the following definitions:
Physics Simulation (PS):
$P_{\theta}(X)\rightarrow X^{\prime}.$ (1)
Physics simulation is the process evolving the input observation $X$ into the output observation $X^{\prime}$ using the physical model $P_{\theta}$, where the observations $X,X^{\prime}$ can be from different simulation time steps.
Physics Understanding (PU):
$X\rightarrow P_{\theta}.$ (2)
Physics understanding is the process inferring the underlying physical model $P_{\theta}$ from the observation $X$ such as video data. Physics understanding can also be conducted to only infer the physical parameters $\theta$ given a predefined physical model $P$.
Generation (G):
$G(X)\rightarrow X^{\prime}.$ (3)
Generation is the process creating new content $X^{\prime}$ from the input condition $X$ using the generative model $G$. The input and output can take various forms of modalities depending on the specific task. If the generation process is general-purposed and does not necessarily involve a strong understanding of the physical world, we call it as physics-unaware generation (PUG).
2.2 Formulation of Physics-Aware Generation
Based on the above concepts, we further define the concept of physics-aware generation (PAG) as the generation process with a strong understanding of real-world physics. We divide physics-aware generation into two major categories, i.e., physics-aware generation with explicit physical simulation (PAG-E), and physics-aware generation without explicit physical simulation (PAG-I). This division is based on whether the generative model utilizes a physical simulation model explicitly to improve physical awareness. Table I provides a comparison of these concepts.
Within the category of physics-aware generation with explicit physical simulation (PAG-E), we further identify the following common paradigms of how the physical simulation is incorporated into the generative model.
Generation to Simulation (GtS): This paradigm is a sequential composition where the simulation process follows the generation process:
$P_{\theta}(G(X))\rightarrow X^{\prime}.$ (4)
Simulation in Generation (SiG): In this paradigm, the simulation model is incorporated as a part or a sub-module of the generation model:
$G_{P_{\theta}}(X)\rightarrow X^{\prime}.$ (5)
Generation and Simulation (GnS): In this paradigm, the generation and the simulation happen simultaneously, such as by using a shared model $M$:
$M_{P_{\theta},G}(X)\rightarrow X^{\prime}.$ (6)
Simulation-Constrained Generation (ScG): In this paradigm, the simulation model is used to provide constraints or knowledge to the generation model:
$G(X)\rightarrow X^{\prime}\;\;subject\;to\;\;P_{\theta}(X)\rightarrow X^{\prime}.$ (7)
Generation-Constrained Simulation (GcS): In this paradigm, the generation model is used to provide constraints or knowledge for the simulation model:
$P_{\theta}(X)\rightarrow X^{\prime}\;\;subject\;to\;\;G(X)\rightarrow X^{\prime}.$ (8)
Simulation-evaluated Generation (SeG): In this paradigm, physical simulation is used to evaluate the generative model or the generated content is intended for deployment in simulation environments.
These paradigms are illustrated in Fig. 2. In the following sections, we will cover physical simulation in Section 3, physical understanding in Section 4, physical-aware generation in Section 5, and physical evaluation in Section 6.
Figure 2: Paradigms of incorporating explicit physical simulation with generative models.
Figure 3: Examples of physical materials, simulation methods, and physical engines used in physics-aware generation. Images of physical materials are generated by GPT-4o.
3 Physical Simulation
Physical simulation refers to the use of simulated environments or models that mimic real-world physical systems, allowing the generative models to learn and infer without direct interaction with the physical world. In the following, we summarize the core elements of physical simulations commonly used in physics-aware generation research, along the dimensions of physical material properties, simulation methods, and off-the-shelf physical engines.
3.1 Physical Materials
In this section, we introduce several physical materials widely used in physical simulation when integrated with physics-aware generation, as detailed below.
Rigid Body. This refers to a material or object that does not deform when subjected to applied forces [83, 84, 85, 86, 87, 88, 89, 90, 91]. In physics-aware generative models, rigid bodies are often used to represent components that maintain their shape and volume. These models simulate rigid body dynamics, including translation, rotation, and collision, to predict the movement of objects.
Soft Body. Soft body is a material [92, 85] that can undergo large deformations when subjected to external forces such as biological tissues, rubber, gels, cloth, and polymers. They exhibit properties like elasticity, viscoelasticity, and non-linear behavior, and they often have complex internal structures.
Newtonian Fluids. Newtonian fluids [93, 89, 94, 95] are fluids whose viscosity (resistance to flow) remains constant regardless of the applied stress or the rate of flow. Common examples include water, air, and most gases. In the modeling process, fluid motion can be described using standard equations like the Navier-Stokes equations.
Non-Newtonian Fluids. These are fluids whose viscosity changes when a stress is applied or when the rate of shear changes [93, 96, 94]. Examples include ketchup, cornstarch in water, and blood. Non-Newtonian fluids can exhibit behaviors such as shear-thinning (viscosity decreases with increasing shear rate), shear-thickening (viscosity increases with increasing shear rate), or even more complex behaviors like viscoelasticity.
Plasticine. Plasticine [93, 88, 86, 97, 98, 94, 91], as a soft and viscoelastic material, is often modeled using viscoelastic and soft body physics approaches. These methods capture both the elastic (rebound) and viscous (flowing or slow deformation) aspects of plasticine’s behavior, making it suitable for simulations of materials that deform and retain new shapes.
Elastic Entities. Elastic materials [94, 88, 98, 99, 93, 84, 100, 86, 96, 97, 101, 91] are characterized by their ability to return to their original shape after deformation, following a linear or nonlinear stress-strain relationship.
Granular Media. Granular media [91, 96], such as sand [94, 102, 103, 93] and snow, behave in complex ways that combine both solid and fluid properties. Their behavior is governed by discrete particle interactions, which are often modeled using Discrete Element Method (DEM) or adapted continuum models that capture phenomena like flow, compaction, and jamming.
Metals. Metals [96, 104, 102] have the ability to undergo elastic and plastic deformations. They are described using elasticity and plasticity models, including work-hardening and yield criteria. These models are crucial for simulating metal behavior under different loads.
Solid. Solids [87, 88, 89] are essential components in physical models due to their definite shape, strength, and mechanical properties. Whether it is modeling stress and strain, plasticity, or thermal expansion, understanding the behavior of solid materials is crucial for accurate simulations and predictions.
Cloth Materials. Cloth [90, 89] exhibits unique behavior due to its flexibility, anisotropy, and resistance to deformation. In simulations, it is usually modeled using techniques like mass-spring systems or particle-based systems to simulate deformations and interactions with forces like wind and gravity.
Articulated Bodies. Articulated bodies [105, 104, 95] are central to many physical models, especially in systems with connected, movable parts like robots, human bodies, and mechanical systems. Their kinematic and dynamic behavior is complex, accounting for motion, forces, collisions, and constraints. These models are used in fields where understanding the interaction of interconnected bodies is crucial.
Thin-shell Objects. Thin-shell objects [106, 105] in physical models refer to structures that have a small thickness compared to their other dimensions (length and width) but still exhibit significant mechanical behavior that needs to be modeled accurately.
3.2 Physical Parameters
Physical materials are substances that have specific properties, such as density, elasticity, and thermal conductivity, which define their behavior under different conditions [101, 107, 97, 108, 95, 83, 109, 110, 99, 94, 111, 86, 105, 92, 93, 88, 89, 87, 102, 98, 104]. These properties are quantified by physical parameters, which are measurable values that describe the material’s response to external forces, heat, or other stimuli. For example, the Young’s modulus of a material defines its stiffness, while thermal conductivity measures its ability to conduct heat. The relationship between materials and their physical parameters is crucial for predicting how materials will behave. The definition of physical parameters is usually correlated with the selection of the physical material and the simulation method. We provide some examples of common parameters for each physical material as shown in Figure 3.
3.3 Simulation Methods
Below we introduce methods for simulating and analyzing physical dynamics, which appear in the literature of physics-aware generation reviewed in this survey.
Continuum Mechanics-based Method (CMBM). Continuum mechanics-based method [106] solves the governing partial differential equations that describe the behavior of materials as continuous media, focusing on stress, strain, and motion in solids, fluids, and gases. These simulations use numerical methods like FEM or Computational Fluid Dynamics to approximate solutions for complex, real-world systems involving large deformations, fluid flow, and heat transfer.
Material Point Method (MPM). The MPM [103, 91, 96, 86, 88, 97, 84, 101, 98, 93, 94, 107, 112] is a hybrid simulation technique that combines Lagrangian particles with an Eulerian grid to model the behavior of materials. In this method, particles carry material properties such as mass, velocity, and stress, while a background grid is used to solve the governing equations of motion. The particles interact with the grid, allowing MPM to efficiently handle large deformations, phase transitions, and multiphase flow while capturing both material behavior and complex interactions.
Finite Element Method (FEM). FEM [110, 96, 103, 100] discretizes a physical domain into small, interconnected elements and solves the governing equations of motion or energy using approximation functions within each element. These element-level equations are then assembled into a global system, which is solved to obtain the behavior of the entire system under various loads and boundary conditions.
Position-Based Dynamics (PBD). PBD [91, 102] is a physics simulation method that focuses on solving constraints directly in terms of particle positions, rather than velocities or forces. The technique iteratively adjusts particle positions to satisfy physical constraints, such as distance, volume preservation, or collision, ensuring stability and efficiency, especially for real-time applications involving deformable objects and soft materials.
Eulerian Method (EM). EM [93, 112, 94, 95] is a grid-based method where the physical domain is fixed, and quantities like velocity, pressure, or temperature are computed at fixed points in space. As the material flows through the grid, EM solves the governing equations (e.g., conservation of mass, momentum) at each grid point to model the evolution of physical fields over time.
Lagrangian Method (LM). LM [93, 112, 100, 109, 94, 95] follows individual particles or material points as they move through space and time, solving the equations of motion based on their positions, velocities, and forces. In practice simulation, LM is particularly useful for tracking the deformation and motion of materials with moving boundaries or interfaces, such as in fluid dynamics or solid mechanics.
3D Spring-Mass Model. This is a physical simulation technique where a deformable object is discretized into masses connected by springs [99], each spring representing elastic forces that govern the object’s deformation. The motion of the masses is computed over time by solving the equations of motion using numerical methods, accounting for forces such as spring tension, damping, and external interactions.
Newtonian Dynamics. The Newtonian Dynamics represents a physical simulation method [103] that models the motion of objects based on Newton’s laws of motion, where the force acting on an object is equal to its mass times its acceleration. It uses numerical integration techniques to solve the equations of motion over time, simulating the object’s response to forces such as gravity, collisions, and applied external forces.
$\Phi$-Flow. This is a physics-based simulation method that uses a level-set function $\Phi$ to represent fluid interfaces and track free-surface flows or multiphase interactions [95]. The method solves the Navier-Stokes equations for fluid dynamics while evolving the level-set function to handle complex interfaces and simulate realistic fluid behaviors.
3.4 Physics Engines and Platforms
To perform physics simulations, several popular engines and platforms have been developed, as summarized below.
Bullet Physics. This is an open-source physics engine widely used in games, animation, and robotics simulations [113]. It enables to support rigid body dynamics, soft body dynamics, collision detection, and physical constraints.
Havok Physics. Havok [114] is a commercial physics engine widely used in games and real-time simulations. It offers efficient rigid body, fluid, cloth, and vehicle physics simulations.
NVIDIA PhysX. This is a physics engine developed by NVIDIA that supports GPU-accelerated real-time simulation [115]. It includes particle systems, cloth, rigid body dynamics, and fluid simulations, optimized with high-performance.
Unity Physics. The built-in physics engine of Unity [91, 116] supports both 2D and 3D physics simulations. It provides collision detection, rigid body dynamics, cloth, and soft body physics.
Unreal Engine Physics (Chaos Physics). The physics engine in Unreal [90] supports large-scale destruction and fluid/cloth simulations, also providing realistic rigid body and soft body physics.
Open Dynamics Engine. This an open-source physics engine [117] focusing on rigid body dynamics and collision detection, widely used in games, robotics, and physical simulation systems.
Box2D. Box2D [118] is a lightweight, open-source 2D physics engine designed specifically for 2D simulation, offering efficient rigid body dynamics and collision detection.
Mantaflow. Mantaflow [119] is an open-source fluid and smoke simulation engine, with the capability to support high-quality fluid, fire, and gas simulations.
Simulink/Matlab (Simscape Multibody). This is a platform within Matlab used for multi-body dynamics and mechanical system simulations [120]. It is highly suitable for detailed modeling and simulation of mechanical systems, robotics, and vehicles.
Blender. This is a free and open-source 3D creation suite [103, 89] that supports modeling, sculpting, animation, rendering, and simulation. It offers a powerful and versatile toolset for productions and scientific simulations.
Isaac Gym. Isaac Gym [109] is a high-performance physics simulation platform developed by NVIDIA, designed specifically for training and testing robotic systems using reinforcement learning. It provides a GPU-accelerated framework that allows to simulate multiple environments in parallel, enabling faster training of agents.
Vortex Studio. This is a high-performance simulation platform primarily designed for robotics, vehicles, and mechanical systems [121]. Its most attractive feature is its ability to offer real-time physics simulation, with a focus on large systems and complex dynamics.
PyBullet. PyBullet is a Python interface for Bullet Physics [122, 123, 90, 108], commonly used in robotics and reinforcement learning, providing simulations of rigid body dynamics in machine learning environments.
Gazebo. Gazebo is a robot simulation platform that provides high-fidelity physics and sensor modeling [124]. It integrates with Robot Operating System for realistic robotics simulations.
Genesis. Genesis [125] is an open-source platform for embodied artificial intelligence that provides a high-fidelity physics environment to simulate robots and their interactions with the physical world. It focuses on helping agents learn through physical interaction with environments, supporting tasks like manipulation, locomotion, and real-world decision-making.
Taichi. Taichi [103] is a platform combining the Taichi programming framework to enhance the efficiency and accuracy of physical system simulations. By integrating optimization and learning algorithms, it enables adaptive modeling, real-time simulation adjustments, and data-driven insights for robotics, virtual environments, and scientific simulation.
4 Physical Understanding
Physical understanding in machine learning aims to infer underlying physical models, laws, or parameters from observational data such as images and videos. This capability is crucial for tasks that require modeling real-world dynamics, including object motion, material behavior, and environmental interactions. While extensive research has been dedicated to physical reasoning and model discovery[126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155], this area is largely orthogonal to physics-aware generative modeling, which is thus out of the core focus of this survey. Nonetheless, accurately estimating physical parameters remains a foundational component in enabling physically consistent generation. In the context of physics-aware generative AI, determining the physical parameters that drive simulations is essential for producing realistic and physically plausible outputs. We identify three major approaches to obtain physical parameters as below.
Manually-Set Physical Parameters. In this approach, domain experts explicitly define the physical properties and constraints used in simulations[100, 96, 91, 123, 103, 112, 108, 110, 156, 109, 83, 95, 104, 111, 105]. These parameters may include material properties (e.g., mass, friction, elasticity), environmental conditions (e.g., gravity, fluid dynamics), and initial configurations. While this method is practical and widely used, it often lacks scalability and adaptability to complex or diverse scenarios.
Automatically-Learned Physical Parameters. Data-driven models can automatically infer physical parameters by learning from visual observations [93, 94, 99, 101, 98, 97, 92, 85] to avoid inflexible pre-defined parameters. These parameters can be estimated either separately in a standalone stage or optimized jointly with the parameters of the generative model.
LLM-Reasoned Physical Parameters. Recent multimodal large language models (LLMs) enable reasoning about physical systems using both textual and visual information [86, 107, 84, 102, 90, 157, 89, 87, 88, 106]. By leveraging contextual knowledge and commonsense reasoning, LLMs prompted with descriptions of objects can infer their physical materials and plausible physical configurations for simulation.
Method Year Paradigm G-Mo G-Md P-Mo P-Pa P-S
[…]
TABLE II: Summary of Physics-Aware Generation Methods with Explicit Physical Simulation (PAG-E).
[…]
5 Physics-Aware Generation
In this section, we first provide an overview of the primary classes of generative models employed in physics-aware generation. We then introduce the main paradigms for incorporating physics-based constraints and knowledge into these generative models.
5.1 Generative Models
We begin by introducing widely used models including Generative Adversarial Networks (GANs), Diffusion Models (DMs), and neural rendering methods—including Neural Radiance Field (NeRF) and Gaussian Splatting (GS)—which, although traditionally viewed as rendering or reconstruction frameworks, are interpreted as generative models in a broader sense in this survey. These generative models can handle a variety of data modalities (e.g., images, videos, 3D, 4D, or interactive environments) and encompass various domains (e.g., general data, human-centric data, and indoor scenes).
5.2 Generative Adversarial Networks
A generative adversarial network[3] consists of two neural networks—a generator and a discriminator—engaged in a competitive process. The generator seeks to create synthetic data that closely resembles real-world data, while the discriminator attempts to distinguish between real and generated data. This adversarial dynamic drives both networks to improve iteratively, resulting in a generator capable of producing highly realistic data samples. Specifically, the generator $G$ and the discriminator $D$ play a two-player minimax game with the following loss function:
[…]
where $x$ is a real sample from the data distribution and $G(z)$ is a sample generated using the latent code $z$. StyleGAN[158] can provide an automatically learned, unsupervised separation of high-level attributes in latent codes, and later versions further fix characteristic artifacts in high-resolution generation[159]. GAN has been the dominant generative model before the popularity of diffusion models.
5.3 Diffusion Models
Diffusion models are a class of generative models that have gained significant attention in recent years for their ability to produce high-quality synthetic data[77]. Compared to GANs, DMs are known for their stability during training but lower efficiency during sampling[5, 6]. Diffusion models leverage a sequential process of transforming simple noise distributions into complex data distributions through a series of learned denoising steps. Diffusion models simulate a gradual corruption of data by adding Gaussian noise over multiple steps in a forward process:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\boldsymbol{I}),$ (10)
where $t$ is the diffusion step, $x_{t}$ is the noisy data at step $t$, $\beta$ is a schedule parameter. This process is then reversed using a neural network parameterized to denoise the data step-by-step:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\boldsymbol{\mu}_{\theta}(x_{t},t),\sigma_{t}^{2}\boldsymbol{I}),$ (11)
where $\sigma$ is the variance schedule. This denoising process can be parameterized in several ways, including mean prediction, noise prediction, and clean data prediction[76]. This iterative refinement allows diffusion models to generate detailed and diverse outputs while maintaining strong theoretical foundations. Score-based models are of an equivalent formulation as diffusion models but from the differential equation perspective[160, 161, 162]. The denoising process can be accelerated by samplers such as DDIM[4] and DPM-Solvers[163]. Latent diffusion models[164] conduct the forward and reverse processes in a high-level latent space and delegate low-level reconstruction to an auxiliary VAE. Techniques like classifier guidance and classifier-free guidance further improve controllability and generation quality[165]. Diffusion model and its later development[166, 167, 168, 169, 170, 171, 172] have been greatly advanced generative tasks for visual data.
5.4 Neural Radiance Field
Neural radiance field[7] is an approach to scene representation and novel view synthesis, which leverages neural networks as implicit representation to model the volumetric properties of a scene. NeRF utilizes a multi-layer perceptron (MLP) to encode a continuous volumetric scene as a function mapping 3D spatial coordinates $x,y,z$ and viewing directions $\theta,\phi$ to radiance density $\sigma$ and color $c=(r,g,b)$ values:
$c,\sigma=F(x,y,z,\theta,\phi).$ (12)
By optimizing the network to minimize the error in rendering the observed views of a scene, NeRF can synthesize photorealistic novel views by querying the MLP along camera rays and using volume rendering techniques to project the output colors and densities into an image. PixelNeRF[173] enables neural scene representation from only one single observed image. MIP-NeRF[174] reduces aliasing artifacts and significantly improves the ability to represent fine details. DNeRF[175] extends neural radiance fields to dynamic objects with motions. These advancements have propelled the neural radiance field into widespread applications in physics-informed scene representations for dynamic and large-scale environments.
5.5 Gaussian Splatting
Gaussian splatting[8] is an approach to scene reconstruction and novel view synthesis that represents a scene as a set of dense, overlapping Gaussian “blobs” in 3D space with learned parameters including the mean $\mu_{i}$, covariance $\Sigma_{i}$, opacity $\alpha_{i}$, and view-dependent color $c_{i}(\theta,\phi)$:
$\mathcal{G}(x,y,z,\mu_{i},\Sigma_{i},\alpha_{i})\cdot c_{i}(\theta,\phi).$ (13)
Instead of relying on volumetric grids or implicit neural fields, Gaussian splatting is an explicit radiance field-based scene representation. During rendering, the Gaussians are “splatted” onto the image plane, with their contributions composited together to form a final image. By jointly optimizing the positions, shapes, and colors of the Gaussians, Gaussian splatting achieves photorealistic renderings with compelling speed and accuracy, making it an attractive alternative or complement to grid-based and neural implicit methods like NeRF. 4D Gaussian Splatting (4D-GS)[176] goes beyond static scenes and extends to dynamic scenes as a 4D representation. SplatterImage[177] maps a single input image to one 3D Gaussian per pixel, enabling fast scene reconstruction from a single view. LGM[178] is proposed for 3D generation, which generates 3D Gaussians from multi-view images synthesized by external models at inference time from text or single-image inputs
5.6 Physics-Aware Generation w/ Explicit Simulation
In the following, we review papers working on physics-aware generation with explicit physical simulation (PAG-E), which are grouped into six categories according to the paradigm of how physics simulation (Sim) is integrated with the generative model (Gen). Note that one paper can usually couple multiple paradigms, but is classified based on the most evident one.
5.6.1 Paradigm 1: Gen-to-Sim (GtS)
This category of papers usually appends physical properties to a generative representation in a post-processing manner to make it simulatable and interactable.
Simulation Elements in NeRF Field. PIE-NeRF[100] uses Poisson disk sampling to distribute “particles" in the NeRF density field, which form simulation elements through Voronoi grouping. Quadratic generalized moving least square (Q-GMLS) strategy and Lagrangian dynamics are then applied to the elements for simulation, enabling users to interact with the NeRF scene using external forces. Video2Game[90] automatically converts a single video of a real-world scene into an interactive virtual environment. To achieve this, a NeRF-generated scene is segmented into individual objects. Each segmented object is then assigned physical parameters such as mass, friction, and collision geometry (e.g., sphere, box, convex polygon) to be simulated with rigid-body physics in real time using WebGL-based game engine.
Gaussian Blobs as Simulation Elements. PhysGaussian[96] first builds a Gaussian Splatting with anisotropic regularization, and then imparts physically grounded behavior to the Gaussian kernels through Material Point Method (MPM). The method applies continuum mechanics to evolve 3D Gaussian kernels, where the kernels are treated as discrete particle clouds for discretizing the continuum. This allows the system to model deformation realistically by tracking physical quantities like stress and strain. Similarly, GASP[103] also leverages the Material Point Method (MPM) to simulate physical behaviors. GASP converts Gaussian components into triangle-based meshes, on which the MPM is applied, and the results are re-parameterized back into Gaussian components for rendering. Spring-Gau[99] first reconstructs static Gaussian, from which anchor points are sampled and are connected via “springs" to model elastic behavior. The model uses differentiable simulation to optimize learnable stiffness and damping parameters from video observations, to transform the Gaussian representation to be dynamic and simulatable. Phy124[112] generates physics-consistent 4D content from a single image, by evolving dynamic 3D content adhering to natural physical laws. Specifically, 3D Gaussians are generated from a single image using diffusion prior and then animated by attaching an MPM simulator.
Physical Feature Field. Feature Splatting[84] extends Gaussian Splatting by embedding semantic features extracted from vision-language models. Such feature-carrying 3D Gaussians bridge the gap between static 3D scene representations and dynamic physical behaviors. The physics engine simulates physical interactions by treating Gaussian centroids as particles assigned with material properties identified by natural language semantics. Phys4DGen[107] utilizes the Segment Anything model and large language models to infer the material composition and physical properties of different object components in a scene represented by Gaussian Splatting. This allows for more accurate physical simulation by assigning realistic material behaviors to different object parts. SimAnything[86] employs a multi-modal large language model to predict the mean physical properties at the object level, which further helps the estimation of the probability distribution of physical properties at a particle level for MPM simulation.
Applications in VR and Robotics. VR-GS[91] introduces a framework designed for interactive and physically realistic manipulation of 3D content in Virtual Reality (VR), which integrates an XPBD physics simulation model with 3D Gaussian Splatting to support real-time and physics-aware interaction. LIVE-GS[102] further incorporates GPT-4 to analyze and infer the physical properties of objects directly from images, bypassing the manual tuning in VR-GS. DreMa[123] combines scene reconstruction with physics-based simulation to create a manipulable, object-centric world model for robotics, enabling robots to imagine new object configurations and predict the outcomes of their actions.
5.6.2 Paradigm 2: Sim-in-Gen (SiG)
This paradigm features physics simulation integrated directly into the generative model, functioning as a core sub-module.
Conditions from Simulation. In GPT4Motion[89], the Blender-based physics simulation is integrated into the video generation pipeline. Specifically, GPT-4 translates user prompts into Python scripts that control Blender’s physics engine to simulate realistic motion. The output of Blender—rendered as edge and depth maps—is then fed into ControlNet-modified Stable Diffusion to generate videos that are both visually consistent and physically plausible. MotionCraft[95] enhances pretrained image diffusion models and introduces motions by warping the noise latent space with optical flow derived from physics simulation. This process allows the model to generate temporally consistent frames that evolve according to physically accurate dynamics such as fluid dynamics, rigid body motion, and multi-agent interactions. Similarly, by using image-based warping with simulated motion dynamics, PhysGen[87] generates physically plausible video outputs given an input image and user-defined forces and torques. The simulation is based on rigid-body dynamics governed by Newton’s Laws, using physical parameters inferred by large foundation models.
Simulation as Optimization. PhyCAGE[156] treats MPM simulation as a physics-aware optimizer by evolving the particle system based on physical laws. The gradient of the loss function is interpreted as the initial velocity of the particles derived from 3D Gaussians, and this velocity is passed into the MPM simulation to optimize the system over sub-steps before each gradient descent step. PhysDiff[85] generates physically plausible human motions by integrating physics-based constraints directly into the sampling process of diffusion models. At each denoising step, the intermediate motion is passed through a physics simulator for correction. The physically-corrected motion then feeds back into the sampling process and guides subsequent denoising steps, steering the generation toward physically plausible motions.
Generative Programs for Simulation. AutoVFX[106] works on photorealistic and physically plausible video editing through natural language instructions, which combines LLM-based code generation and physics-based simulation. User instructions are transformed into executable code, enabling edits such as object insertion, material changes, dynamic interactions, and particle effects. The code is executed in the Blender engine to render the scene with photorealistic lighting and materials to produce the edited video. Similarly, GPT4Motion[89] uses large language models to automate the generation of Blender Python scripts that drive physical simulations, integrating semantic understanding with procedural generation to create complex and physics-based animations to guide video generation.
5.6.3 Paradigm 3: Gen-and-Sim (GnS)
This paradigm involves simultaneous or interconnected operation of generation and simulation processes with tight coupling.
Joint Geometry and Physics Learning. PAC-NeRF[93] addresses the challenge of inferring both the geometric and physical parameters of objects solely from multi-view video data, using a hybrid Eulerian-Lagrangian representation of the scene. Concretely, the Eulerian grid representation is used for NeRF to learn the geometry, while the Lagrangian particle representation is used for simulation to learn the physical parameters. iPAC-NeRF[94] further proposes Lagrangian Particle Optimization to directly optimize the positions and features of particles in the Lagrangian space, leading to dynamic refinement of the geometric structure across the entire video sequence while adhering to physical constraints.
Alternating Simulation and Generation. PhysMotion[88] follows a generation-simulation-generation process for image-to-video generation. Firstly, the foreground object in the input image is converted into a coarse 3D Gaussian Splatting representation. Then, the model undergoes simulation using MPM, applying physical laws to simulate how the object would behave under forces, to produce a coarse video depicting physics-grounded dynamics. Finally, the coarse video is further refined using diffusion-based video enhancement to improve visual realism.
5.6.4 Paradigm 4: Sim-Constrained Gen (ScG)
This is a paradigm where the simulation imposes constraints or guidance on the training of generative models to improve physical awareness.
Loss Functions from Simulation. PhysComp[110] creates 3D models from single images while ensuring physical compatibility. It is constrained by a simulation-based physical model—specifically, the static equilibrium constraint—which ensures the generated 3D shapes behave realistically under physical forces. PhyRecon[109] incorporates simulation as loss functions to constrain and guide the generation process toward physically plausible 3D scenes. This is achieved through the integration of a differentiable particle-based physical simulator, which directly influences the neural implicit surface representation to improve stability and model physical uncertainty. Atlas3D[83] generates 3D models that are self-supporting, incorporating a standability loss to penalize rotational instability to ensure the generated model maintains its upright orientation during simulation, and a stable equilibrium loss to encourage resilience against minor perturbations. Mezghanni et al.[108] introduce two differentiable physical loss functions, i.e., a connectivity loss to regularize that the generated 3D shapes are single connected components, and a stability loss to promote physical stability under gravity.
Simulation for Data Filtering DiffuseBot[92] employs simulation as a component for data filtering during its embedding optimization phase. Specifically, 3D robot models generated from diffusion models are evaluated using a differentiable physics simulation to assess the performance of each robot based on task-specific metrics. A filtering mechanism then selects the top-performing designs to be retained, gradually skewing the sampling distribution of the generative model toward successful robot designs.
Reinforcement Learning with Physical Feedback. Furuta et al.[157] addresses the challenge of generating realistic dynamic object interactions in text-to-video models, by using reinforcement learning (RL) fine-tuning from external feedback—particularly from vision-language models (VLMs). The VLM-based feedback mechanism serves as a form of simulation or evaluation of physical realism, which constrains and guides the generation model toward producing physically plausible videos.
5.6.5 Paradigm 5: Gen-Constrained Sim (GcS)
In this paradigm, the generation model acts as a guidance or prior knowledge for the simulation process.
Score Distillation Sampling. Physics3D[97] combines elastoplastic and viscoelastic MPM with a video diffusion model. To optimize physical parameters, it utilizes Score Distillation Sampling (SDS) from a pretrained video diffusion model. The diffusion model provides motion pattern priors learned from video data, guiding the simulation toward physically plausible behaviors without requiring manual specification of material properties. DreamPhysics[98] further proposes Motion Distillation Sampling (MDS) to better capture motion-specific priors and reduce color bias in the optimization.
Learning Physical Parameters from Generated Data. PhysDreamer[101] synthesizes realistic interactive dynamics for objects by extracting physical dynamics priors from video generation models. Different from score distillation sampling, this method learns such priors from generated videos. The method first uses a pre-trained image-to-video generation model to create a reference video of a static object undergoing plausible motions. Then it optimizes physical parameters such as Young’s modulus in MPM simulation by maximizing the visual similarity between the simulation-rendered video and the generated reference video.
5.6.6 Paradigm 6: Sim-evaluated Gen (SeG)
In this paradigm, generation is intended for simulation-based deployment, focusing on the utility in simulated environments.
Liu et al.[104] generates physically plausible articulated 3D assets from limited data. By using auxiliary penetrations penalization and collision-based shape optimization, this method produces articulated meshes that behave correctly under simulation. PhysPart[111] generates 3D replacement parts that fit precisely and move smoothly, essential for applications such as 3D printing and robotic manipulation. The approach integrates geometric conditioning through classifier-free guidance and enforces physical constraints via stability and mobility losses during the sampling process. PhyScene[105] generates high-quality interactive 3D scenes tailored for embodied artificial intelligence, which leverages a conditional diffusion model to capture scene layouts and employs physics-based guidance functions, ensuring the scene is physically functional and suitable for embodied agent interactions.
5.7 Physics-Aware Generation w/o Explicit Simulation
In this section, we introduce representative works in generative AI for vision that exhibit a degree of physical awareness without relying explicitly on physical simulation (PAG-I).
Physical Awareness in Large Video Models. Recent advancements in large-scale video generation models, such as OpenAI’s Sora[20], OpenSora[179], CogVideoX[180], VideoCrafter[29], LAVIE[35], ModelScope[181], demonstrate emergent physical reasoning capabilities. These models, scaled up and trained on vast datasets of internet-scale video content, have shown promising abilities to implicitly capture and replicate certain physical dynamics and causal relationships present in the real world[182]. This emergent physical understanding enables them to generate complex scenes with coherent object interactions, realistic motions, and plausible temporal dynamics. This progress positions them as potential foundations for universal simulators capable of modeling diverse physical phenomena, despite lacking explicitly programmed physics engines[45]. However, this implicit physical reasoning remains in its early stage and systematic evaluations are still open areas. For example, PhyGenBench[183] introduces a comprehensive benchmark designed to evaluate the physical commonsense understanding of text-to-video generation models. The findings highlight that current models, despite their impressive visual generation capabilities, struggle to represent basic physical laws accurately. Kang et al.[184] investigate whether large-scale video generation models can autonomously learn and generalize physical laws solely from visual data. It is found that scaling model size and dataset do not help with out-of-distribution physical generalization. Current models primarily rely on the existence of closely similar training examples rather than abstracting general physical rules, underscoring the need for more targeted methods to develop fully physics-aware generative models.
Physical Awareness from Large Language Models. Large language models can also supply physical knowledge for improving the generation of visual data. PhyT2V[185] addresses the lack of physical realism in text-to-video models by integrating large language models to refine textual prompts iteratively. The process first extracts objects and relevant physical rules from the initial text prompt, and then uses a video captioning model to generate semantic descriptions of the generated video and compare it with the original prompt to identify mismatches. Finally, the text prompt is refined based on the extracted physical properties to correct the mismatches.
Generative Interactive Dynamics. Generative interactive dynamics focus on modeling the natural dynamics of objects without requiring explicit physical simulations, implicitly learning the physical relationships and interactions. Blattmann et al.[186] makes an early attempt to predict object dynamics in response to localized user interactions within an image-to-video formulation. Generative Image Dynamics[187] generates realistic scene motion from a single image, via a latent diffusion model that predicts spectral volumes representing dense long-term pixel trajectories. Motion-controlled video generation and video editing belong to another form of generative dynamics[188, 57]. Motion Prompting[56] leverages motion trajectories as conditioning signals, allowing for fine-grained and flexible control of local and global scene motions. Similarly, VideoComposer[43] and Yoda[55] utilize motion vectors as explicit temporal control signals to allow for precise capture of inter-frame dynamics and patterns. Motion Dreamer[189] decouples motion reasoning from high-fidelity video generation, using intermediate motion representations such as optical flow, instance segmentation maps, and depth maps. Motion Guidance[190] controls the diffusion generation process using gradients derived from optical flow estimators, enabling detailed motion edits without model retraining.
Physical Domain Data Generation. CoCoGen [191] models domain-specific data in physics, such as Darcy Flow and Burgers Equation, using score-based generative models. The method enforces physical consistency by injecting discretized partial differential equation information directly into the sampling process of score-based models. Cao et al.[192] enhance video diffusion models with latent physical knowledge for generating videos that require high physical accuracy, such as fluid simulation and typhoon phenomenon videos. The method pretrains masked autoencoders on physical phenomena data to capture latent physical knowledge, which is then projected into a quaternion space to construct pseudo-language prompts that guide the video diffusion model in creating physically plausible outputs.
6 Physics Evaluation
Most of the image and video generation methods are evaluated based on FID [193], CLIPSIM [194], CLIP-FID [195], and FVD [196] to evaluate visual quality and diversity.
However, there are a number of gaps in these metrics. Frechet Inception distance (FID) [193], Clip Similarity (CLIPSIM) [194], CLIP-FID [195] and Inception score (IS) [193] are more biased toward the alignment between images’ contents and text prompt. For video generation, the detection of unrealistic motions violating physical laws is more important, which cannot be detected by these metrics effectively. Furthermore, the evaluation of FVD requires a reference video dataset, which is difficult to have for newly generated scenes. These challenges prevent these metrics from effectively evaluating physical common sense. Some works use Vision Language Models (VLMs) as evaluators [197, 198], yet these works only evaluate the spatial relationships. Most of them face challenges when generalizing to physical correctness.
In this section, we introduce the benchmarks and metrics designed for evaluating the physics awareness of image or video generation models.
6.1 Benchmarks
Category Physical laws
Mechanics gravity, buoyancy, solid pressure, atmospheric pressure,
elasticity, friction, surface tension
Optics reflection, refraction, scattering, dispersion,
interference and diffraction, straight-line propagation
Thermal Solidification, Melting, Liquefaction, Boiling,
deposition, Sublimation
Material color, hardness, solubility, combustibility and flame reaction,
acidity, redox potential, dehydrating properties
TABLE III: Taxonomy of physical commonsense in PhyGenBench[199].
PhyBench. PhyBench[200] and PhyGenBench[199] define physical commonsense as four types, including mechanics, optics, thermal and material properties associated with physical laws as in Table III. PhyBench [200] constructs the benchmark by extending the designed implicit prompt describing a type of physical commonsense. For example, given “pressure" topic, the authors design a prompt “an empty plastic bottle at the bottom of the sea". This is because the bottle would stand a very large pressure when staying under the water. After that, “empty plastic bottle" will be replaced by other objects, and the scenario "at the bottom of the sea" is also replaced by other scenarios to describe other topics. This replacement process is conducted using GPT-4.
PhyGenBench. Due to the limited scope of PhyBench, PhyGenBench [199] constructs the benchmark with 4 categories from Table III for video generation through five steps:
  • Conceptualization: Identify key physical commonsense from four types of physical commonsense
  • Prompt Engineering: Craft initial prompts to describe the underlying physical phenomenon.
  • Prompt Augmentation: Adding additional details to prompt
  • Diversity Enhancement: Use GPT-4o to do object substitution on the augmented prompts
  • Quality control: Review the prompts and associated physical laws for accuracy and relevance.
VideoPhy. Besides the four physical categories defined in Table III, VideoPhy[201] categorizes the physical interaction types into three categories which are Solid-Solid, Solid-Fluid and Fluid-Fluid. An example of Solid-Solid is the prompt “A cat is playing with a bottle," where the interaction between “cat" and “bottle" is solid-solid, or “Water flows in an open pipeline" is the interaction between “water" and “pipeline" describing Solid-Fluid. Similarly, the prompt “Rain is dropping onto a river" is a Fluid-Fluid interaction. Based on these definitions, VideoPhy [201] constructs a total of 688 captions for the benchmark.
6.2 Evaluation metrics
There are two main approaches for the evaluation metrics of physical awareness. The first approach is based on human evaluation, where humans will give the score to the image or video in terms of physical commonsense with defined criteria. The second way is automatic evaluation.
For human evaluation, this method achieves very high expectations for the evaluation process [200, 199, 201, 202]. However, this results in exorbitant cost due to the involvement of humans, which is also slower.
For automatic evaluation, there are two approaches, i.e., obtaining scores or yes/no answers from LVMs/LLMs models [200, 199, 201] or obtaining specialized quantitative scores.
Specialized Quantitative Scores. In [192], the authors used Stream Function Error (SFE) [203, 204] for assessing streamline accuracy, Smoothness Error (SE) [205] for measuring stable propagation characteristics. Gradient Smoothness (GS) and Continuity Score (CS) [204] are used to evaluate the reasonableness and consistency of the phenomenon without using ground truth. Q-Criterion Error (QCE) [206] and Vorticity Error (VE) [207] are used to measure the preservation of vortex structure and vorticity accuracy accordingly.
LVMs-Based Scores. There are two approaches by which we can utilize the LVMs model to evaluate the physical commonsense. The first method is to ask the LVM model to give a score for scene and physical correctness assessments directly according to some criteria [200]. The model is also required to provide justification for its score. In [199], the author further improves the evaluation by evaluating three different scores, which are “Key Physical Phenomena Detection", “Physics Order Verification," and “Overall Naturalness Evaluation". The scores obtained from these three criteria will be averaged to achieve the final score.
  • Key Physical Phenomena Detection: This score is to decide if the physical phenomena occur in the video.
  • Physics Order Verification: This score will assess if the physical phenomena occurred in the correct order.
  • Overall Naturalness Evaluation: This score aims to assess the overall naturalness of the video.
Besides from asking the score directly from the LVMs, [201] asks the LVM to answer if the generated video/image has satisfied Semantic Adherence (SA) or Physical Commonsense (PC). The answer from LVM is required to be “Yes" or “No". The final score will be calculated as:
$s_{\theta}(\tau_{t}(x))=\frac{p_{\theta}(\text{Yes}|\tau_{t}(x))}{p_{\theta}(\text{Yes}|\tau_{t}(x))+p_{\theta}(\text{No}|\tau_{t}(x))}$ (14)
where $p_{\theta}(\text{Yes}|\tau_{t}(x))$ is the probability that the LVM model will answer “Yes" with $t\in\{SA,PC\}$.
7 Discussions
In this section, we discuss related concepts and provide comparisons with physics-aware generation.
Semantic Understanding vs. Physical Understanding. Semantic understanding and physics understanding serve distinct but complementary roles in how models interpret and generate visual data. Semantic understanding involves interpreting and recognizing concepts and objects within visual data, focusing on identifying “what" and “where" are present in a scene, by learning static knowledge mapping from pixels to latent features. In contrast, physics understanding pertains to the ability to simulate dynamic interactions and transitions governed by physical laws, such as motion, collisions, forces, and material behaviors, addressing the “how" and “why" of visual data. In this sense, physics understanding is characterized by its dynamic and predictive nature, which explains why temporal modeling of videos is especially important in this domain. Physics-aware generation aims to combine both understandings to produce content that is both semantically meaningful and physically accurate.
Geometry Understanding vs. Physical Understanding. Geometry understanding involves perceiving and modeling the extrinsic structural properties of objects and scenes. This includes aspects such as shape, size, pose, depth, and positions[208, 209, 210, 211, 212]. Physics understanding focuses on modeling the intrinsic physical properties and behaviors of objects and environments. This involves reasoning about how objects move, deform, and interact under physical laws[93]. In this sense, physics understanding is characterized by its causality and interaction modeling. Integrating both geometry and physics understanding is essential for creating embodied models that not only perceive but also understand and interact with the physical world.
8 Future Directions
In this section, we discuss several potential future directions along the promise of physics-aware generation.
Better Evaluation of Physics-Awareness. Although standardized benchmarks tailored to physical awareness have been introduced recently[183], the evaluation remains ad-hoc or largely relies on manual assessment. The evaluation of physical consistency and dynamic realism in generative models remains a significant challenge that hinders progress. One potential direction is to integrate generative models with established physics engines to evaluate physical plausibility, by converting and deploying generated content within simulation platforms. Incorporating task-oriented evaluations can also be beneficial to assessing physical reasoning capabilities. For example, assessing how well generated content performs in downstream tasks—such as object manipulation in robotics or environmental modeling in autonomous driving—can offer feedback on a model’s utility and physical validity. It is also promising to explore the alignment between generated content and the physical world using embodied agents as an evaluation measurement, for example, by using the behavior similarity of embodied agents running in the simulated and the real world.
Explainability by Physics-Awareness. The integration of physics-aware mechanisms into models holds immense potential for enhancing model explainability. Incorporating explicit physical laws into generative models can offer interpretable pathways linking inputs to outputs, enabling users to trace how specific forces, constraints, and interactions govern the evolution of generated content. By mapping model decisions to well-understood physical principles, the generation process becomes more transparent and aligned with human intuition. Combining physics-aware generative models with existing explainable AI methods, such as saliency maps[213] and feature attribution[214], can further enhance interpretability. For example, visualizing how physical parameters influence specific regions of generated outputs can reveal the model’s reliance on physical properties during generation.
Physics-Augmented Foundational Models. It is promising to augment large foundational models with physical understanding and utilize them for physics-aware generation[215][216]. This includes a versatile understanding of physical world by unifying various perception foundation models[217, 218, 219, 220, 221, 222, 223, 224], and enhanced reasoning abilities from large language models augmented with 3D[225, 226] and physical knowledge[227]. Another future direction is physics-guided pretraining and representation learning, by leveraging large-scale synthetic datasets and differentiable physics integration. It is also interesting to explore self-supervised physical learning, by implementing learning objectives that encourage the discovery of physical laws in large models.
Neural-Symbolic Hybrid Models. The integration of neural and symbolic reasoning presents a promising avenue for advancing physics-aware generation, where symbolic systems can offer rigorous reasoning based on structured knowledge of physics and causality[228, 229]. Embedding symbolic physical constraints, such as differentiable temporal logic[230], can guide generative models to produce outputs that are physically valid and temporally consistent. Symbolic representations also excel in compositional reasoning, allowing systems to understand complex entities and relationships. By incorporating symbolic graphs or ontologies, generative models can achieve more complex compositional generation.
Generative Simulation Engine. Generative simulation engine offers the potential to generate physical simulation directly from textual or semantic prompts, with Genesis[125] as a recent example. Such text-to-simulation models aim to convert high-level textual descriptions into fully interactive, physics-consistent virtual environments. For instance, a command such as “simulate a landslide on a steep mountain after heavy rain" could trigger the generation of a realistic 3D terrain, modeled with appropriate soil and fluid dynamics, and simulate the resulting landslide using physics-based solvers. Generative simulation engines can automate and speed up the design of physical simulation, which can be further integrated into larger systems for physics-aware generation in various modalities.
Robotics and Embodied AI. A natural application of physics-aware generation is to be used as world models or simulators to train robots and embodied agents. By generating physically realistic synthetic data that accurately reflects real-world dynamics, robots can be trained more effectively in simulation and seamlessly be transited to real-world operations[231, 62, 232]. It is also interesting to explicitly inject physical reasoning capabilities into Vision-Language-Action models[233, 226] to better predict actions and outcomes in complex and transferred environments with physical knowledge that can better generalize.
Applications in Interdisciplinary Domains. Physics-aware generative AI also holds transformative potential across many application domains. Physics-aware generative models can improve climate modeling by enhancing the fidelity, resolution, and scale of simulations[234]. In healthcare, physics-aware generation can significantly enhance surgical training and planning, such as building surgical simulations that accurately replicate the physical properties of tissues to provide realistic training environments[235, 236].
9 Conclusion
This survey comprehensively reviewed the rapidly evolving field of physics-aware generative AI in vision, highlighting the efforts made to enhance the physical realism and functionality of generated visual content. The core of this survey focused on the integration of physical simulation into generative models, delineating between physics-aware generation with explicit simulation and without explicit simulation. Physics-aware generation stands at a transformative juncture, bridging the gap between virtual and physical realities. We hope that this survey serves as a helpful resource and inspiration for future research in the field of physics-aware generation.